On June 30, 2018, John and Co. issued ten year, 6% bonds with face value of $500,000. The bonds are dated December 31, 2017. Interest will be paid annually each December 31. The bonds were sold to yield 5%. John and Co. Follows IFRS. Provide the journal entries for the bond issue and show calculations.
Answer #1

ENTRY FOR ISSUE OF BONDS IN THE BOOKS OF JOHN & CO.

PARTICULARS DEBIT CREDIT

BANK A/C DEBIT 525000

TO 6% BONDS 500000

TO PREMIUM ON BONDS 25000

(BEING BONDS ISSUED AT 5% PREMIUM)

CALCULATIONS:

PREMIUM ON BONDS=FACE VALUE*50%

500000*5%=$25,000

AMOUNT IN BANK WOULD BE PREMIUM+FACE VALUE=500000+25000=$525000

CALCULATION OF INTEREST AMOUNT ON DATE 31 DEC 2017:

AMOUNT OF INTEREST= FACE VALUE*BOND RATE

=500000*6%

=$30,000

NOTE: BONDS ARE ISSUED ON 1 JUNE 2018 BUT ON BONDS IT IS DATED BY 31 DEC 2017 SO INTEREST FOR THE WHOLE YEAR IS TO BE PAID.

ENTRY FOR INTEREST PAID:

INTEREST ON BONDS A/C DEBIT 30000

TO BANK A/C 30000

(BEING INTEREST PAID FOR WHOLE YEAR AT 6%)